OpenStudy (anonymous):

How do you solve this!?!?!? 2pi(x/2pi)=5pi

OpenStudy (jdoe0001):

well \(\pi \) is really just a constant

OpenStudy (anonymous):

yes but how is the answer to x 5pi

OpenStudy (anonymous):

x=5pi

OpenStudy (jdoe0001):

\(\bf \cancel{ 2\pi }\left(\cfrac{x}{\cancel{ 2\pi }}\right)=5\pi\)

OpenStudy (anonymous):

oh i see. i feel stupid

OpenStudy (anonymous):

i have another question

OpenStudy (imstuck):

Don't ever say you feel stupid...THAT'S a stupid thing to say!

OpenStudy (anonymous):

hahaha Here's my other question: Part 4: What is the coordinate point associated with this radian measure?

OpenStudy (jdoe0001):

with "this" radian measure?

OpenStudy (jdoe0001):

ohh you mean \(\large 5\pi\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

well, a full circle revolution is \(2\pi\) 2 revolutions are \(4\pi\) so... where would you end up with \(4\pi + \pi\quad ?\)

OpenStudy (anonymous):

5pi?

OpenStudy (jdoe0001):

well.... yes